OpenStudy (anonymous):

what is the value of u in terms of x in the equation ((3x+y)/(2x))=15/2

OpenStudy (anonymous):

you are back @gohangoku58

OpenStudy (anonymous):

yes :) and you want 'y' in terms of x, not u, right ?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

Yes i want the value of y sorry typo

OpenStudy (anonymous):

so multiply 2x on both sides first ?

OpenStudy (anonymous):

This is a proporsion

OpenStudy (anonymous):

oh wait never mind i can cross multiply

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes! cross multiplying is bette :)

OpenStudy (anonymous):

what you got after cross-multiplying ?

OpenStudy (anonymous):

You need to cancel the y or the x out i think the x

OpenStudy (anonymous):

as of now, nothing will get cancelled. did you try cross-multiplying ?

OpenStudy (anonymous):

Nvr mnd

OpenStudy (anonymous):

thanks anyways

OpenStudy (anonymous):

2(3x+y) = 15* 2x 6x+2y = 30 x 2y = 24x y= 12x welcome anyways :)